Minimize Swarm Operational News

Swarm Alpha: EFI LP and LPI data gap on 18 December 2018

02 January 2019

Electric Field Instrument (EFI) LP and LPI data measured by Swarm Alpha are not available on 18 December 2018. The reason for this data gap is still under investigation.

Swarm Alpha: VFM data gap from 18 to 19 December 2018

02 January 2019

Due to a science packet corruption which occurred on the VFM (Vector Field Magnetometer) on board Swarm Alpha, the VFM science data (at 50 Hz rate) were not received from 18 December 2018 at around 18:13 UTC to 19 December 2018 at 16:46 UTC when the VFM was restarted. As a consequence, no L1B VFM data are available for this time window.

Swarm Bravo: Gap in EFI data from 16 to 17 October 2018

19 December 2018

We inform the Swarm users that an issue was registered on-board Swarm Bravo during the EFI-TII Gain Correction Map application. This issue caused a data gap in EFI data from 16 October 2018 13:33:57 UTC to 17 October 2018 08:28:17 UTC

Swarm Alpha VFM issue from 29 April to 01 May 2017: data regenerated

20 November 2018

With reference to the news published on 04 May 2017, we inform the Swarm users that the Swarm Alpha MAGx_LR_1B data from 29 April to 01 May 2017 have been regenerated and now can be found in the ESA FTP server.

Swarm Bravo VFM issue from 11 to 13 September 2017: data generated

20 November 2018

With reference to the news published on 14 September 2017, we inform the Swarm users that the Swarm Bravo MAGx_LR_1B data from 11 to 13 September 2017 have been generated and now can be found in the ESA FTP server.

Swarm Charlie STR issue from 07 to 25 May 2018: data regenerated

20 November 2018

With reference to the news published on 29 May 2018, we inform the Swarm users that the issue present in the STR data, caused by 1-pixel shift in STR CHU-3, has been fixed. All the Swarm Charlie Magnetic data from 07 to 25 May 2018 have been regenerated and now can be found in the ESA FTP server.

Extension of Swarm 2 Hz TII Cross-Track data flow dataset

08 November 2018

We would like to inform the Swarm users that the 2 Hz TII Cross-Track data flow dataset has been extended in the ESA FTP server, including additional data during 2017 for spacecraft A, and sparse days in 2014, for A and B spacecraft.

Swarm Alpha: ACC data gap from 16 to 17 October 2018

22 October 2018

Due to an anomaly that occurred on Accelerometer on-board Swarm Alpha, the instrument was switched off from 16 October 2018 00:53:21 UTC to 17 October 2018 07:05:06 UTC.

Changes in the Swarm Level 1B and Level 2 Cat-2 operational processors

18 September 2018

On 17 September 2018 an improved version for both L1B and L2 Cat-2 Operational Processors was transferred into operations. We are in the process of publishing the entire datasets generated with the latest processor versions (L1B and L2 Cat-2 outputs with the following Product Baseline and File Counter: ORBATT 0401; MAGNET 0505; PLASMA 0501; ACCELE 0401; IBI 0301 ; FAC 0301; TEC 0301)

Swarm Bravo: VFM data gap from 19 to 27 August 2018

29 August 2018

Due to a science packet corruption which occurred on the VFM (Vector Field Magnetometer) on board Swarm Bravo, the VFM science data (at 50 Hz rate) were not received from approximately 10:00 UTC on 19 August to 07:10 UTC on 27 August 2018 when the VFM was restarted.

Swarm Charlie: Gap in EFI data on 07 August 2018

16 August 2018

A gap in data acquired by the Electric Field Instrument (EFI) on the Swarm Charlie spacecraft occurred on 07 August 2018. The duration of this gap was from 14:46:31 to 14:53:12 UTC and was due to a gain map update.

Swarm Alpha: L1B ASM data gap from 24 to 26 July 2018

06 August 2018

Due to a test performed on the ASM (Absolute Scalar Magnetometer) commanded in burst mode on board Swarm Alpha, the magnetic field intensity derived from the ASM in the MAGA_LR_1B and MAGA_CA_1B products is set to zero from 24 July 2018 at 14:44 UTC to 26 July 2018 at 07:00 UTC.

Unexpected Swarm A GPS receiver reboot on 26 July 2018

31 July 2018

The Swarm A GPS receiver rebooted on 26 July 2018 at 05:07:02 UTC autonomously. The GPS receiver data and time synchronization were available again at 05:29:47 UTC. The satellite's orbit and attitude seem unaffected by the anomaly. While the investigation is ongoing, it seems that there is no impact on the good health of the GPS receiver.

Swarm satellite thermo-optical properties and external geometry published

22 June 2018

Modelling of non-gravitational forces such as drag or solar radiation pressure requires knowledge of the thermo-optical properties and geometry of the outer surfaces of the spacecraft. These are now provided in a memorandum and can be accessed here.

Swarm data access simplified

21 June 2018

All validated Swarm Level 1b and Level 2 products are now freely accessible to all users via anonymous FTP or any HTTP browser.

Earth Observation data distribution services improvement

14 June 2018

ESA has started a process to improve the Earth Observation data distribution services, aiming at facilitating access to data and information for the end users, initially by shortening the access path and reviewing authentication and authorisation processes.

Swarm Alpha: VFM data gap from 02 to 03 June 2018

05 June 2018

Due to an anomaly that occurred on the VFM (Vector Field Magnetometer) on board Swarm Alpha, the VFM science data (at 50 Hz rate) was not received from 02 June 2018 at 09:52 UTC to 03 June 2018 at 09:14:34 UTC.

Swarm Charlie: Telemetry data gap on 30 May 2018

01 June 2018

On 30 May 2018 a technical issue caused a telemetry data gap in S/C C from 21:14:47 to 23:12:23 UTC.

Swarm Charlie STR data issue fixed on 25 May 2018

29 May 2018

With the reference to the news published on 25 May 2018, we would like to inform the users that after  performing a power cycle on the Swarm Charlie CHU3 on 25 May 2018 at 12:44 UTC, the anomalous variation in the IBAs is no more present.

Swarm Charlie STR data issue from 07 May 2018

25 May 2018

An anomalous variation in the Inter Boresight Angle (IBA) for Camera Head Unit (CHU) pairs A-C and B-C has been observed for Swarm Charlie from 07 May 2018 at 09:34 UTC.

Showing 1 - 20 of 169 results.
of 9